OpenStudy (anonymous):

find the sum of 1/4+2/7

OpenStudy (andriod09):

Hey deejay12, welcome to openstudy! The first thing to do, is to change the denominators so that their the same. What number would you multiply 4, and 7 by that would make them the same denominator?

OpenStudy (andriod09):

Nfgallimore, please remember that giving another just an answer is against the OS CoC which you agreed to when you signed up. Feel free to look here for more info: http://www.openstudy.com/code-of-conduct

OpenStudy (anonymous):

4 * 7 = ? the common denominator would be 7/? + 8/? 7 * 4 = ?, 1 * 7 = 7